Please solve these question:Which of the following is an irrational number.a.19 b.2.5 c.3.333 d.5\7
Answer by greenestamps(13200)



None of those numbers is irrational.

19 = 19/1
2.5 = 5/2
3.333 = 3333/1000
5/7 = 5/7